Decide whether or not the following quantified propositions are true or false, where
in each case the universal set is the set of positive integers N+. Justify each of your conclusions
with a proof or a counterexample.
(a) ∀n(2
n ≥ n
2
)
(b) ¬∃n(n
2 = 15)
1
Expert's answer
2019-11-11T10:37:30-0500

(a) We should decide whether for all positive integers n, the inequality 2n>=n2 holds.


The statement is false. For n=3 we have that 23=8<9=32.


But for all positive integers "n\neq 3" this inequality holds.

For n=1 we have 2>1, for n=2 we have 22=22. To prove that the inequality holds for n>=4 we will use mathematical induction. The base case is n=4, 24=42. Let us prove the inductive step. Suppose that 2n>=n2. Then 2n+1=2"\times" 2n>2n2 by induction hypothesis. Since n>=4, we have that n2>2n+1. Hence, 2n2>n2+2n+1=(n+1)2.Thus, 2n+1>(n+1)2. Since both the base case and the inductive step have been performed, by mathematical induction the inequality holds for all n>=4.


(b) The function "n\mapsto n^2" is monotone. We have that 32=9, 42=16. Thus, it does not exist positive integer n such that n2=15. The statement is true.